OpenStudy (anonymous):

How to solve this 3 equations

OpenStudy (anonymous):

I mean any tricks to solve

OpenStudy (johnweldon1993):

The second equation....that is indeed yz + yz right?

OpenStudy (anonymous):

wait do we add those

OpenStudy (anonymous):

\[\huge y ^{2}+yz+yx=-12\]

OpenStudy (anonymous):

\[\huge x ^{2}+xy+xz=18\]

OpenStudy (anonymous):

\[\huge z ^{2}+zx+zy=30\]

ganeshie8 (ganeshie8):

yeah smells like (x+y+z)^2 but how are ou gona find x,y,z values ? are you supposed to solve them in integers ?

OpenStudy (anonymous):

yeah the question is silent on that

OpenStudy (anonymous):

By adding we get:- \[\huge (x ^{2}+y ^{2}+z ^{2})+2(yz+zx+yx)=36\]

OpenStudy (anonymous):

is this some kind of expansion

OpenStudy (phi):

you can factor each into e.g. y(x+y+z)= -12 if you divide them you can get ratios, and put e.g. y and z in terms of x

OpenStudy (anonymous):

\[\huge x+y+z=6\]

OpenStudy (anonymous):

@phi i didn't get u

OpenStudy (anonymous):

But how is x+y+z=6 helping

OpenStudy (anonymous):

can it also be -6

ganeshie8 (ganeshie8):

\(\pm 6\)

OpenStudy (phi):

example: x/y = -3/2 so y= -⅔ x x/z = 3/5 so z= 5/3 x use those in x(x+y+z)= 18

OpenStudy (anonymous):

yeah

OpenStudy (johnweldon1993):

Yeah I believe there are 2 separate solutions for this

OpenStudy (anonymous):

yeah there are 2 sollutions given

OpenStudy (phi):

factor y out of the first equation x out of the second z out of the third what do you get ?

ganeshie8 (ganeshie8):

or x(x+y+z) = 18 pluging the two x+y+z valyes to get the two solutions for x

OpenStudy (anonymous):

cool!

OpenStudy (anonymous):

Nice application of the formula

OpenStudy (anonymous):

i mean identity

OpenStudy (anonymous):

THANKS ALL!!

ganeshie8 (ganeshie8):

i think 3 independent equations and 3 unknowns cannot have infinitely many solutions ! i was thinking there might be infiniftely many real solutions earlier which is just wrong as wolfram confirms it : http://www.wolframalpha.com/input/?i=solve+y%5E2%2Byz%2Byx+%3D+-12%2C+x%5E2%2Bxy%2Bxz%3D18%2Cz%5E2%2Bzx%2Bzy%3D30

OpenStudy (anonymous):

3 equations for 3 unknowns ,

OpenStudy (phi):

just to finish this approach example: x/y = -3/2 so y= -⅔ x x/z = 3/5 so z= 5/3 x use those in x(x+y+z)= 18 x(-⅔ x + 5/3 x)= 18 2x^2= 18 x= ± 3

ganeshie8 (ganeshie8):

they are like 3 surfaces intersecting one another exactly at two points or something

OpenStudy (anonymous):

my brain was first thinking of finding out x in terms of x,y,z and then substitute that chunk of shi.t into the second equation Xd

OpenStudy (anonymous):

I have learnt something that whenever there are some big looking equations , there is always a hidden trick to solve them

OpenStudy (phi):

and y= -⅔ x gives y= ± 2 (but the -2 goes with the +3 x solution) z= 5/3 x gives z=± 5 (3,-2,5) and (-3,2,-5)

OpenStudy (anonymous):

yes they r same differ in sign though i got them